OpenStudy (anonymous):

Solve for x, given the equation Square root of x minus 5 + 7 = 11.

OpenStudy (anonymous):

\[\sqrt{x-5}+7=11\] subtract \(7\) to get \[\sqrt{x-5}=4\] then square to get \[x-5=16\]

OpenStudy (anonymous):

is it extraneous?

OpenStudy (anonymous):

it is x−5−−−−−√+7=11 subtract 7 to get x−5−−−−−√=4 then square to get x−5=16

OpenStudy (anonymous):

what grade are you in veronicafresta